# Repository-level Context Code Translation Benchmark Targeting Rust

### v5

@@ Abstract.
Recent advancements in large language models (LLMs) have demonstrated impressive capabilities in code translation, typically evaluated using benchmarks like CodeTransOcean. However, these benchmarks fail to capture real-world complexities by focusing primarily on simple function-level translations and overlooking repository-level context (e.g., dependencies). Moreover, LLMs’ effectiveness in translating to newer, low-resource languages like Rust remains largely underexplored.
@@
To address this gap, we introduce RustRepoTrans, the first repository-level context code translation benchmark, comprising 375 tasks translating into Rust from C, Java, and Python. Using this benchmark, we evaluate four state-of-the-art LLMs, analyzing their errors to assess limitations in complex translation scenarios. Among them, Claude-3.5 performs best with 43.5% Pass@1, excelling in both basic functionality and additional translation abilities, such as noise robustness and syntactical difference identification. However, even Claude-3.5 experiences a 30.8% performance drop (Pass@1 from 74.3% to 43.5%) when handling repository-level context compared to previous benchmarks without such context. We also find that LLMs struggle with language differences in complex tasks, and dependencies further increase translation difficulty.
copyright: acmlicensedjournalyear: 2025
1. Introduction
@@
Code translation, or code migration, is the process of converting a software project from one programming language to another (Pan et al., 2023), often driven by the need to adapt to different runtime environments, improve performance, or enhance security (nee, 2018, 2020b, 2020c, 2020d, 2020a; Thönes, 2015). The rise of languages like Rust (rus, 2024) and Cangjie (can, 2025) has increased demand for translation. However, migrating legacy code remains challenging due to semantic mismatches, differences in standard libraries, and the need to maintain functional equivalence (Mateus et al., 2023; Kontogiannis et al., 2010).
@@
Large language models (LLMs) have shown promise in translating widely used languages like Java and Python (Pan et al., 2024, 2023), but their effectiveness in translating to newer, lower-resource languages like Rust has not been thoroughly explored. Rust, an emerging language known for its safety guarantees, is increasingly in demand for translating projects from other languages to enhance memory safety and reliability (Security, 2024). However, its distinct features and limited training data pose challenges for automated translation, potentially limiting existing models’ performance (rus, 2024).
Existing code translation benchmarks, while covering Rust to some extent, have several limitations in capturing real-world scenarios. Most of them focus on function-level translation pairs (Roziere et al., 2020; Lu et al., 2021; Zheng et al., 2023; Jiao et al., 2023), often translating self-contained functions with minimal dependencies (Puri et al., 2021; Yan et al., 2023; Zhu et al., 2022a). Even file-level benchmarks only provide limited context for making isolated functions executable. Furthermore, many datasets are sourced from online coding platforms or artificially created examples, which differ greatly from real-world development practices (Puri et al., 2021; Yan et al., 2023; Zhu et al., 2022a; Roziere et al., 2020; Lu et al., 2021; Zheng et al., 2023; Jiao et al., 2023). In contrast, real-world translation occurs at the repository level, where developers incrementally translate code while preserving the project’s architecture. This involves managing complex dependencies, cross-file references, and inconsistencies between large-scale projects, aspects that current benchmarks fail to capture.
@@
To bridge this gap, we introduce RustRepoTrans, the first benchmark to evaluate code translation with repository-level context, specifically targeting Rust. Unlike function-level benchmarks that assess translation in isolation, RustRepoTrans captures the complexities of real-world migration by incorporating repository-level context, including dependencies, cross-file interactions, and architectural constraints. These elements are crucial for ensuring that translated code integrates seamlessly into existing projects rather than functioning as isolated snippets. Our benchmark comprises 375 curated translation tasks derived from real-world projects spanning C, Java, and Python to Rust, providing a more faithful assessment of LLMs’ translation capabilities. While we focus on Rust due to its growing adoption in systems programming and the inherent challenges of translating into it, our methodology is generalizable to other programming languages, paving the way for broader repository-level evaluation.
Building on RustRepoTrans, we assess state-of-the-art LLMs in translating code with complex repository-level context, analyzing challenges beyond existing benchmarks. We identify key factors affecting task difficulty and common errors, while also evaluating noise robustness, syntactical difference handling, and code simplicity during translation.
@@
Our experiments show that LLMs struggle with repository-level code translation, with compilation errors reaching 94.8%, exposing the gap between current evaluations and real-world performance. Claude-3.5 performs best with 43.5% Pass@1, excelling in basic functionality and additional translation abilities like noise robustness and syntactical difference identification. However, even Claude-3.5 sees a 30.8% drop (Pass@1 from 74.3% to 43.5%) when handling repository-level context. Dependency-related errors, including function and variable resolution issues, account for 61.9% of failures, highlighting the challenge of interconnected code. We also find that code complexity correlates with translation effectiveness. These insights reveal LLMs’ limitations in complex code migration, guiding future improvements in model design and evaluation.
@@
The contributions of this paper are as follows. Data and code are publicly available at  (rep, 2024).
@@
  • The first repository-level code translation benchmark, RustRepoTrans, with 375 tasks from C, Java, and Python to Rust, is introduced for realistic evaluation.
@@
  • Four state-of-the-art LLMs are evaluated on this benchmark, assessing their performance in real-world translation tasks.
@@
  • LLM errors are categorized into 10 types, revealing limitations, especially in handling dependencies.
2. Related Work
2.1. Code Translation
Code translation enhances software interoperability and facilitates legacy system updates, boosting productivity and reducing manual costs (Lu et al., 2021; Puri et al., 2021). Early methods relied on rule-based systems with handcrafted syntactic and statistical machine translation (SMT) models, improving accuracy by mapping frequent code patterns (Nguyen et al., 2013; Chen et al., 2018). Neural models, including RNNs, improved learning from data but struggled with complex syntax and idioms (Dong and Lapata, 2016; Yin and Neubig, 2017, 2018).
Despite advancements with LLMs, translation errors and adaptability issues persist (Zhu et al., 2022a; Rithy et al., 2022). Jiao et al. (Jiao et al., 2023) highlighted LLMs’ difficulty with complex translations, while Pan et al. (Pan et al., 2024) showed that context-rich prompts improve reliability by 5.5%. Yang et al. (Yang et al., 2024) and Macedo et al. (Macedo et al., 2024a) further refined prompt engineering for accuracy and consistency.
Our work introduces RustRepoTrans, which enables a more realistic evaluation of LLMs by incorporating repository-level dependencies and context. This approach allows us to better assess existing code translation techniques in handling complex code translation tasks.
2.2. Code Translation Benchmarks
Recent advances in neural code translation models have been fueled by high-quality, large-scale parallel datasets. However, existing datasets exhibit significant limitations in programming language coverage, scale, and balance, particularly for Rust. Many datasets focus narrowly on a few languages, such as the early studies between Java and C (Nguyen et al., 2013) and more recent work like CodeTrans (Lu et al., 2021) and AVATAR (Ahmad et al., 2021), which primarily cover Java and Python.
To address these gaps, some datasets, such as CoST (Zhu et al., 2022b) and CodeNet (Puri et al., 2021), have aimed to enhance language diversity and quality. However, issues like erroneous solutions and limited support for Rust persist. The latest benchmark, CodeTransOcean (Yan et al., 2023) and XCodeEval (Khan et al., 2024a), expands coverage to low-resource languages but still lacks adequate support and error analysis for Rust, hindering comprehensive research in this area.
Furthermore, many existing benchmarks rely on data from question-and-answer websites, which do not reflect real-world software development contexts. These datasets fail to capture the complex dependencies and architectural considerations inherent in actual projects, where code translation often requires maintaining the original architecture and addressing intricate cross-file dependencies. This disconnect underscores the need for more representative benchmarks like RustRepoTrans, which aim to provide a more realistic assessment of code translation capabilities.
RustRepoTrans is the first repository-level context code translation benchmark focusing on Rust, designed to account for complex dependencies and derived from real-world projects, distinguishing it from previous benchmarks.
3. RustRepoTrans Benchmark
@@
3.1. Benchmark Format
Figure 1. RustRepoTrans Format Example
@@ 3.1. Benchmark Format
3.2. Benchmark Construction Method
@@
The construction process is divided into two parts: Functionally Equivalent Code Pairs Extraction and Dependency Extraction. This approach allows us to obtain functionally equivalent code pairs along with their corresponding dependencies and test cases from real open-source projects.
@@
Functionally Equivalent Code Pairs Extraction In this part, we focus on extracting functionally equivalent code pairs (source function and target function), with the target function written in Rust, from GitHub projects. The process involves five steps, detailed as follows.
@@
Migration Project Selection. In this step, we select projects that have been rewritten in Rust from other languages, specifically C, Java, and Python, due to their popularity and likelihood of having such rewritten versions. These projects are expected to contain functionally equivalent function pairs between their Rust and original language versions. We identify suitable projects by searching GitHub for terms like “Rust version” and “implemented in Rust”. focusing on larger projects (applying qualifier “size:>1000”) to ensure sufficient function pairs. After locating these projects, we verify their versions through documentation and code review. In this way, we identify pairs of source project and target project.
@@
Functions Pools Extraction. In this step, we extract all functions from a pair of projects (source and target). For the Rust target project, we focus on candidate functions with associated test cases to ensure verifiability. We first identify all test cases and then extract the functions they cover by analyzing the external functions invoked. For the source project, we extract all implemented functions for comprehensive coverage. This extraction uses the static code analysis tool tree-sitter (tree sitter, 2023). As a result, we obtain two sets of functions: the source functions pool and the target functions pool with test cases.
@@
Similarity-based Candidate Function Pair Extraction. In this step, we extract function pairs from the source and target function pools using a similarity-based approach. Developers often translate code at the function level while maintaining a similar structure across languages. Thus, equivalent pairs usually come from files with similar paths and function signatures (Zhong et al., 2010). For example, the Rust function $pbkdf2$ at $src/ecdh.rs$ corresponds to $PBKDF2$ in the C project at $src/ecdh\_support.c$. We use the BM25 algorithm (Trotman et al., 2014) to calculate the similarity for each target function, identifying the top 10 candidate source functions. This produces a list of Rust target functions with their top-10 source function candidates.
@@
LLM-based Equivalent Function Pair Identification. In this step, we identify the most equivalent function for each Rust target function from its top-10 source function candidates using an LLM, leveraging their code implementation and contextual information (such as file paths). LLMs like GPT-4 excel in code understanding (Li et al., 2023). The prompt used for this identification is shown in Fig. 2. We employ GPT-4o due to its effective balance of efficiency and performance. If none of the candidates are functionally equivalent, the LLM is instructed to select "None."
Figure 2. Prompt for Identifying Equivalent Function Pairs
Manual Verification. Due to the limitations of LLMs, such as the potential for hallucinations, we manually verify the functionally equivalent pairs identified by the models to ensure their actual functional equivalence.
@@
We classify dependencies into three categories: invocation function dependencies, invocation variable dependencies, and data type dependencies. To identify these, we perform static analysis on the entire project to extract custom functions, data types, and global variables. For the target function, we gather import statements, call function identifiers, variable dependencies, and data type dependencies. We match in-file dependencies and, using the import statements, match cross-file dependencies to compile a complete list for the function. Due to tree-sitter’s limitations, we manually reviewed the automatically extracted dependencies to correct any errors or omissions, ensuring each function has a complete and accurate set of dependencies.
3.3. Resulting Benchmark
This process constructs RustRepoTrans with 375 repository-level translation tasks. Table 1 compares it with existing benchmarks. Similar in size to prior datasets, each task includes a manually verified ground truth translation and unit tests, achieving over 90% test coverage, ensuring high quality.
RustRepoTrans has two key features that distinguish it from previous benchmarks: 1) it is the first code translation benchmark to consider repository-level dependencies; and 2) it specifically targets code translation to Rust in realistic programming scenarios. Next, we will discuss these features in more detail.
Table 1. Comparison of Different Benchmarks for Code Translation
Dataset Source Task Level #Tokens # Tasks Source Languages With Dependencies? Target Languages Include Rust? Unit Tests Included ? Golden Answer Verified?
CodeXGLUE(Lu et al., 2021) Lucune, POI, JGit, Antlr Function Level 42.3 1,000 Java, C#
XLCOST(Zhu et al., 2022a) G4G Program Level 202 901 C++, Java, C#, PHP, JavaScript, Python, C
TransCoder-test(Roziere et al., 2020) G4G Function Level 107.0 948 C++, Java, Python Partial
HumanEval-X(Zheng et al., 2023) HumanEval Program Level 97.7 164 C++, Java, Go, JavaScript, Python
G-TransEval(Jiao et al., 2023) HumanEval, G4G, .NET samples Function Level 95.1 400 C++, Java, C#, JavaScript, Python
CodeTransOcean(Yan et al., 2023) Rosetta Code Program Level 448.4 2,878* Java, C++, C#, PHP, Python, Go
RustRepoTrans Github Function Level 150.1 375 C, Java, Python

### v6

# RustRepoTrans: Repository-level Context Code Translation Benchmark Targeting Rust

@@ Abstract
Recent advancements in large language models (LLMs) have demonstrated impressive capabilities in code translation, typically evaluated using benchmarks like CodeTransOcean and RepoTransBench. However, dependency-free benchmarks fail to capture real-world complexities by focusing primarily on simple function-level translations and overlooking repository-level context (e.g., dependencies). Full-repository translation benchmarks significantly exceed the current capabilities of existing models, resulting in performance bottlenecks that fail to provide actionable insights for guiding model development. Furthermore, existing benchmarks do not account for the scenario of incrementally translating new or modified modules from the source to the target language, which demands careful handling of repository-level contexts such as dependencies, cross-module references, and architectural divergence. Moreover, LLMs’ effectiveness in translating to newer, low-resource languages like Rust remains largely underexplored.
@@
To address these gaps, we introduce RustRepoTrans, the first repository-level context code translation benchmark targeting incremental translation, comprising 375 tasks translating into Rust from C, Java, and Python. Using this benchmark, we evaluate seven representative LLMs, analyzing their errors to assess limitations in complex translation scenarios. Among them, DeepSeek-R1 performs best with 51.5% Pass@1, excelling in both basic functionality and additional translation abilities, such as noise robustness and syntactical difference identification. However, even DeepSeek-R1 experiences a 22.2% performance drop (Pass@1 from 73.7% to 51.5%) when handling repository-level context compared to previous benchmarks without such context. Meanwhile, we propose a set of more fine-grained evaluation metrics and an enhanced evaluation framework, enabling a more comprehensive analysis of LLMs’ performance in repository-level context code translation tasks to provide fine-grained insights that can effectively inform the development of code translation techniques.
I Introduction
@@
Code translation, or code migration, is the process of converting a software project from one programming language to another [1], often driven by the need to adapt to different runtime environments, improve performance, or enhance security [2, 3, 4, 5, 6, 7]. The rise of languages like Rust [8] and Cangjie [9] has increased demand for translation. However, migrating legacy code remains challenging due to semantic mismatches, differences in standard libraries, and the need to maintain functional equivalence [10, 11].
In real-world projects, translation rarely rewrites an entire repository at once. Developers typically incrementally translate new or modified modules from the source to the target language [12, 13, 14, 15]. Newly translated code must integrate with existing modules, including components migrated earlier or manually refined, to preserve functional correctness and consistency. Incremental workflows also arise when the source project evolves after partial migration, requiring new features to be ported into the target repository. Such repository-level migration demands careful handling of repository-level context such as dependencies, cross-module references, and architectural divergence. It further supports human-in-the-loop workflows, where developers refine each step under non one-to-one code mapping, reflecting the realities of cross-language migration.
@@
While large language models (LLMs) have shown promise in translating widely used languages like Java and Python [16, 1], but their effectiveness in translating to newer, lower-resource languages like Rust has not been thoroughly explored. Rust is increasingly chosen for migration due to its memory safety and reliability benefits [17], yet its strict ownership model and limited training data pose significant challenges for automated translation [8].
Existing benchmarks can be categorized into dependency-free and full repository, however, both categories of datasets only partially capture these requirements. Most focus on function-level translation [18, 19, 20, 21], where self-contained snippets have minimal dependencies, and even file-level benchmarks [22, 23, 24] provide limited repository context. Many benchmarks are also sourced from online platforms or synthetic examples, diverging from real-world development practices [22, 23, 24, 18, 19, 20, 21]. In contrast, practical Rust migration demands repository-level reasoning, dependency management, and incremental integration—factors largely absent from dependency-free benchmarks.
While repository-level translation benchmarks have emerged [25, 26] (e.g., RepoTransBench), fully translating entire repositories remains highly challenging for current LLMs. In RepoTransBench [25], for example, the best model, Claude-3.5, achieves only 7.33% Success@1, highlighting a large performance gap. Moreover, full-repository translation complicates fine-grained evaluation, as errors may result from the model or from differences in target-language libraries, making it difficult to pinpoint their source. Furthermore, existing repository-level benchmarks generally assume ideal conditions, where function dependencies and project architecture are perfectly aligned between source and target languages. In practice, however, differences in language features, optimization goals, and coding idioms often lead to divergent architectures. Consequently, one-to-one mapping between source and target codebases is rare, reflecting realistic context shifts. As illustrated in Fig. 1, the same function may have different dependencies (e.g., built-in vs. custom utilities) or distinct signatures (e.g., parameter-driven vs. global configuration) across languages. These limitations show that current repository-level benchmarks reveal performance gaps but fail to fully assess LLMs’ ability to handle repository-level context, dependency interactions, and incremental translation. Therefore, an intermediate benchmark is needed—one that incorporates repository-level context while remaining tractable for current models. Such a benchmark can bridge the gap between dependency-free function-level translation and full-repository translation, enabling fine-grained evaluation and better guiding the development of robust code translation techniques.
Figure 1: Motivation Examples from RustRepoTrans of Different Architecture Between Source Language (Python) And Target Language (Rust) Version
@@
To bridge this gap, we introduce RustRepoTrans, the first benchmark to evaluate code translation with repository-level context, specifically targeting Rust. Unlike function-level benchmarks that assess translation in isolation, RustRepoTrans captures the complexities of real-world migration by incorporating repository-level context, including dependencies, cross-file interactions, and architectural constraints. These elements are crucial for ensuring that translated code integrates seamlessly into existing projects rather than functioning isolatedly. Compared to full repository benchmarks, RustRepoTrans focuses on incremental translation scenarios which demands the careful handling of non one-to-one code mapping such as different dependencies or distinct function signatures. Overall, RustRepoTrans comprises 375 curated translation tasks derived from real-world projects spanning C, Java, and Python to Rust, providing a more faithful assessment of LLMs’ translation capabilities. While we focus on Rust due to its growing adoption and inherent challenges of translating to it, our methodology is generalizable to other programming languages, paving the way for broader repository-level evaluation.
@@
Building on RustRepoTrans, our experiments of seven representative LLMs show that LLMs struggle with repository-level context code translation, with compilation errors reaching 92.3%, exposing the gap between current evaluations and real-world performance. DeepSeek-R1 performs best with 51.5% Pass@1, excelling in both basic functionality and additional translation abilities like noise robustness and syntactical difference identification. However, even DeepSeek-R1 experiences a 22.2% performance drop (Pass@1 from 73.7% to 51.5%) when handling repository-level context. Dependency-related errors, including function and variable resolution issues, account for 67.6% of failures, highlighting the challenge of interconnected code. Meanwhile, we propose a set of more fine-grained evaluation metrics such as evaluating noise robustness, syntactical difference handling, and code simplicity during translation and an enhanced evaluation framework, enabling a more comprehensive analysis of LLM performance in repository-level context code translation tasks (in RQ4).
@@
The contributions of this paper are as follows. Data and code are publicly available at [27] and [28].
@@
  • The first repository-level context code benchmark, RustRepoTrans, targeting incremental translation with 375 tasks from C, Java, and Python to Rust, is introduced for realistic evaluation;
@@
  • Seven representative LLMs are evaluated, assessing their performance in real-world incremental translation;
@@
  • LLMs’ errors are categorized into 10 types, revealing limitations, especially in handling dependencies;
  • A set of more fine-grained evaluation metrics of code translation and an enhanced evaluation framework.
II Related Work
II-A Code Translation
Code translation promotes software interoperability and legacy system modernization, enhancing productivity and reducing manual effort [19, 22]. Early approaches employed rule-based and machine learning [29, 30, 31, 32, 33] to capture frequent code patterns. Although LLMs have advanced the field, challenges remain in handling translation errors and adapting to complex scenarios [24, 34]. Jiao et al.[21] underscored LLMs’ difficulty with intricate cases, while Pan et al.[16] demonstrated that context-rich prompts improve reliability. Prompt engineering strategies by Yang et al.[35] and Macedo et al.[36] further enhanced translation consistency and accuracy. LLM have also been applied to the translation of entire repository [37, 38, 39]. Nitin et al. [37] leverage C2Rust to convert C into unsafe Rust and then apply an LLM for safer, idiomatic translation. Zhang et al. [38] translate Go to Rust via predefined rules, code partitioning, and localized checks. However, the evaluation data used in these works are not composed of ground-truth code pairs, only the source language version is genuinely available.
We propose RustRepoTrans, a benchmark that incorporates repository-level context and dependencies, enabling more realistic evaluation of LLMs and their capacity to handle complex code translation tasks.
II-B Code Translation Benchmarks
Existing code translation datasets are generally classified into dependency-free, function-level [18, 19, 20, 21] (or single-file [22, 23, 24]) datasets and full repository-level datasets [25, 26]. The former are often mined from Q&A platforms—e.g., CodeTransOcean from Rosetta Code, XCodeEval from GeeksForGeeks—but fail to reflect real-world software development, lacking the architectural and dependency complexities critical for faithful translation across files. In contrast, RepoTransBench [25] introduces a repository-level benchmark with executable tests, while TRANSREPO-BENCH [26] proposes a Skeleton-Guided-Translation framework for Java-to-C# translation, leveraging coarse-grained architectural guidance for holistic evaluation. Nevertheless, full repository translation remains highly challenging; for example, the best model on RepoTransBench achieves only 7.33% Success@1 [25], underscoring current LLM limitations. Additionally, full-repo evaluation hinders fine-grained assessment due to cross-language inconsistencies such as library discrepancies.
RustRepoTrans, the first repo-level context benchmark, bridges the gap between function-level and full-repo translation by offering a repository-level context benchmark with manageable complexity.
III RustRepoTrans Benchmark
@@
III-A Benchmark Format
Figure 2: RustRepoTrans Format Example
@@ III-A Benchmark Format
III-B Benchmark Construction Method
@@
The construction process is divided into two parts: Functionally Equivalent Code Pairs Extraction and Dependency Extraction, enabling us to obtain functionally equivalent code pairs along with their corresponding dependencies and test cases from real open-source projects.
@@
Functionally Equivalent Code Pairs Extraction In this part, we focus on extracting functionally equivalent source–target function pairs. The pipeline consists of five stages:
@@
Migration Project Selection. In this step, we select projects that have been rewritten in Rust from other languages, specifically C, Java, and Python, due to their popularity and likelihood of having such rewritten versions. We identify suitable projects by searching GitHub for terms like “Rust version” and “implemented in Rust”, focusing on larger projects (applying qualifier “size:>1000”) to ensure sufficient function pairs. After locating these projects, we verify their versions through documentation and code review. In this way, we identify pairs of source project and target project.
@@
Functions Pools Extraction. In this step, we extract all functions from a pair of projects (source and target). For the source project, we extract all implemented functions. For the Rust target project, we focus on candidate functions with associated test cases to ensure verifiability. We first identify all test cases and then extract the functions they cover by the static code analysis tool tree-sitter [40]. As a result, we obtain two sets of functions: the source functions pool and the target functions pool with test cases.
@@
Similarity-based Candidate Function Pair Extraction. In this step, we extract function pairs from the source and target function pools using a similarity-based approach. Developers often translate code at the function level while maintaining a similar structure across languages. Thus, equivalent pairs usually come from files with similar paths and function signatures [41]. For example, the Rust function $pbkdf2$ at $src/ecdh.rs$ corresponds to $PBKDF2$ in the C project at src/ecdh_support.csrc/ecdh\_support.c. We use the BM25 algorithm [42] to calculate the similarity for each target function, identifying the top 10 candidate source functions. This produces a list of Rust target functions with their top-10 source function candidates.
@@
LLM-based Equivalent Function Pair Identification. In this step, we identify the most equivalent function for each Rust target function from its top-10 source function candidates using an LLM, leveraging their code implementation and contextual information (such as file paths). LLMs like GPT-4 excel in code understanding [43]. The prompt used for this identification is shown in Fig. 3. We employ GPT-4o due to its effective balance of efficiency and performance. If none of the candidates are functionally equivalent, the LLM is instructed to select "None."
Figure 3: Prompt for Identifying Equivalent Function Pairs
Manual Verification. Due to the limitations of LLMs, such as the potential for hallucinations, two of the authors with 4-6 years of coding experience are involved as participants to manually double check the equivalent of verified function pair by LLMs to ensure actual functional equivalence. Since functionally equivalent project pairs in different programming languages, sharing the same version, have already been identified during the Migration Project Selection phase using project documentation, participants only need to validate whether substantial project architectural refactoring has occurred, which results in the absence of functionally equivalent function pairs or in cases where functions with identical names implement different functionalities across the two projects.
@@
Dependencies are classified into three categories: function dependencies, variable dependencies, and data type dependencies. To identify these, we perform static analysis on the entire project to extract custom functions, data types, and global variables. For the target function, we gather import statements, call function identifiers, variable dependencies, and data type dependencies. We match in-file dependencies and, using the import statements, match cross-file dependencies to compile a complete list for the function. Due to tree-sitter’s limitations, we manually reviewed the automatically extracted dependencies to correct any errors or omissions, ensuring each function has a complete and accurate set of dependencies.
III-C Resulting Benchmark
This process constructs RustRepoTrans with 375 repository-level translation tasks. Table I compares it with existing benchmarks. Similar in size to prior datasets, each task includes a manually verified ground truth translation and unit tests, achieving over 90% test coverage, ensuring high quality.
RustRepoTrans has two key features that distinguish it from previous benchmarks: 1) it is the first repository-level context code translation benchmark and 2) it specifically targets code translation to Rust in realistic programming scenarios. Next, we will discuss these features in more detail.
TABLE I: Comparison of Different Benchmarks for Code Translation
Dataset Source Task Level #Tokens
# Tasks
Source Languages
Average number
of Dependencies
Target Languages
Include Rust?
Unit Tests
Included ?
Golden Answer
Verified?
CodeXGLUE[19] Lucune, POI, JGit, Antlr Function Level 42.3 1,000 Java, C# 0
XLCOST[24] G4G Program Level 202 901 C++, Java, C#, PHP, JavaScript, Python, C 0
TransCoder-test[18] G4G Function Level 107.0 948 C++, Java, Python 0 Partial
HumanEval-X[20] HumanEval Program Level 97.7 164 C++, Java, Go, JavaScript, Python 0
G-TransEval[21]
HumanEval, G4G,
.NET samples
Function Level 95.1 400 C++, Java, C#, JavaScript, Python 0
CodeTransOcean[23] Rosetta Code Program Level 448.4 2,878* Java, C++, C#, PHP, Python, Go 0
RustRepoTrans Github Function Level 150.1 375 C, Java, Python 5.4